OpenStudy (roadjester):

32. An electron has a kinetic energy of 12.0 eV. The electron is incident upon a rectangular barrier of height 20.0 eV and width 1.00 nm. If the electron absorbed all the energy of a photon of green light (with wavelength 546 nm) at the instant it reached the barrier, by what factor would the electron’s probability of tunneling through the barrier increase?

OpenStudy (roadjester):

@douglaswinslowcooper @ybarrap

OpenStudy (roadjester):

The sum of all the probabilities is going to equal 1 so part of the 1 goes to Transmission, the rest goes to Reflection. \[\large R+T=1\] The equation for transmission is defined as \(\LARGE T=e^{-2CL}\) where \(\LARGE C= \dfrac {\sqrt{2m(U-E)}}{h-bar}\). Based on the given information, if I multiply the top and bottom by the speed of light, then the equation now becomes \(\LARGE C= \dfrac {\sqrt{2mc^2(U-E)}}{(h-bar) *c}\). K=12.0 eV L=1.0 nm U=20.0 eV \(\lambda\)=546 nm Using this information, the energy of the photon is \(\large E=hf={hc\over \lambda}\). The energy of the electron is 511 MeV so doing the substitution, C=14.51/nm. If I do the proportionality \(\huge \dfrac{e^{-2CL}}{e^{-2C'L}}=e^{-2L(C-C')}\). I got C'=12.28/nm but that is incorrect. The final (correct) answer is actually 1.2255 X10^10 m ^-1.

OpenStudy (roadjester):

h-bar is planck's constant divided by 2pi.

OpenStudy (roadjester):

@LastDayWork @Mashy

OpenStudy (anonymous):

Sorry. I do not have a clue, except that 1228 is nearly 1226.